# Fleurs-SLU: A Massively Multilingual Benchmark for Spoken Language Understanding

### v2

@@ Abstract
Spoken language understanding (SLU) is indispensable for half of all living languages that lack a formal writing system, since these languages cannot pair automatic speech recognition (ASR) with language models to benefit from language technology. Even if low-resource languages possess a writing system, ASR for these languages remains unreliable due to limited bimodal speech and text training data. Better SLU can strengthen the robustness of massively multilingual ASR by levering language semantics to disambiguate utterances via context or exploiting semantic similarities across languages. However, the evaluation of multilingual SLU remains limited to shallow tasks such as intent classification or language identification. To address this, we present Fleurs-SLU, a multilingual SLU benchmark that encompasses (i) 692 hours of speech for topical utterance classification in 102 languages and (ii) multiple-choice question answering through listening comprehension spanning 944 hours of speech across 92 languages. We extensively evaluate both end-to-end speech classification models and cascaded systems that combine speech-to-text transcription with subsequent classification by large language models on Fleurs-SLU. Our results show that cascaded systems exhibit greater robustness in multilingual SLU tasks, though speech encoders can achieve competitive performance in topical speech classification when appropriately pre-trained. We further find a strong correlation between robust multilingual ASR, effective speech-to-text translation, and strong multilingual SLU, highlighting the mutual benefits between acoustic and semantic speech representations.111The datasets are available on Huggingface: SIB-Fleurs; Belebele-Fleurs.
Fleurs-SLU: A Massively Multilingual Benchmark for
Spoken Language Understanding
Fabian David Schmidt1, Ivan Vulić2, Goran Glavaš1, David Ifeoluwa Adelani3 1 Center For Artificial Intelligence and Data Science, University of Würzburg, Germany 2 Language Technology Lab, University of Cambridge, United Kingdom 3 Mila, McGill University and Canada CIFAR AI Chair {fabian.schmidt, goran.glavas}@uni-wuerzburg.de iv250@cam.ac.uk david.adelani@mila.quebec
@@ 1 Introduction
Only about half of the world’s living languages possess a formal writing system, underscoring the need for massively multilingual speech technology Ethnologue (2017). Spoken language understanding (SLU) is hence a critical feature of inclusive technology for these languages, which cannot rely on combining automatic speech recognition (ASR) with language models to offload semantic speech understanding. However, the datasets currently available for evaluating SLU in truly relevant languages suffer from significant limitations. For instance, the Minds14 benchmark assesses SLU on 14 exclusively high-resource languages for intent classification, a task that often requires only shallow semantic processing, such as detecting specific keywords Gerz et al. (2021). Likewise, while SpeechTaxi focuses on spoken topical classification of Bible verses in 28 diverse languages, SLU of Bible verse arguably does not reflect real-world usage Keller and Glavaš (2024).
@@
Moreover, even when a low-resource language has a writing system, multilingual ASR often struggles to reliably transcribe that language due to the limited availability of bimodal speech and text data. Pre-trained on the self-supervised wav2vec-BERT objective Baevski et al. (2020), these models embed acoustic features that capture phonetic rather than semantic information Choi et al. (2024). Conversely, multilingual speech models that excel in semantic encoding tend to be more robust in ASR tasks by, e.g., leveraging context-related cues or cross-lingual similarities such as cognates or similar syntactic structures (cf. §5.1). This underscores that robust SLU is an instrumental component of genuinely inclusive speech technology.
@@
To address the shortcomings of multilingual SLU evaluation of prior benchmarks, we compile Fleurs-SLU by realigning datasets derived from Flores-200 (Flores). We first filter out silent instances from Fleurs Conneau et al. (2022) and then map the remaining data back to Flores Team et al. (2022). The resulting dataset is then merged with SIB-200 Adelani et al. (2024) and Belebele Bandarkar et al. (2023) to create a topical utterance classification benchmark for 102 languages, and dataset for multiple-choice question answering (QA) on spoken paragraphs for 92 languages, respectively.
@@
Contributions. 1) To the best of our knowledge, we are the first to present a multilingual SLU benchmark that spans over 100 languages and enables end-to-end SLU evaluation of multilingual speech encoders (mSE). SIB-Fleurs supports semantic utterance classification with 692 hours of speech spanning 102 languages, over 70 more additional languages than existing benchmarks. Belebele-Fleurs provides textual multiple-choice QA from spoken paragraphs, totaling 944 hours of speech across 92 languages and exceeding concurrent work by 18 languages Costa-jussà et al. (2024). 2) We extensively benchmark state-of-the-art speech models on SIB-Fleurs and Belebele-Fleurs. Like prior work, we observe that Cascaded Systems (CS) remain more robust than mSE, enabling SLU competitive with text-based language understanding Keller and Glavaš (2024). We are the first to show that mSE pre-trained on language understanding objectives can yield performance competitive to that of CS on speech classification. We further isolate utterance quality as an important factor in multilingual SLU and show that, in zero-shot cross-lingual transfer (zs-xlt), worse utterance quality can substantially deteriorate transfer performance. 3) We empirically demonstrate that strong multilingual SLU coincides with much more robust multilingual ASR and higher-quality speech-to-English-text translation (S2ETT). This suggests that the pre-training of multilingual speech models should take SLU capabilities into account to make massively multilingual ASR more robust.
@@ 2 Related Work
Modern speech representation models are pre-trained with self-supervised objectives that are optionally followed-up by ASR and speech-to-text translation (S2TT) training. Among these models, mHubert employs the self-supervised wav2vec-BERT 2.0 objective on 90K hours of speech across 147 languages by predicting pseudo-labels derived from clustering raw speech features Boito et al. (2024). MMS-1b is also pretrained with wav2vec-BERT 2.0 on an large corpus of 491K hours of speech spanning 1,406 languages Pratap et al. (2024). Whisper-v3 is a Transformer encoder-decoder that has been multi-task pre-trained on multilingual ASR, speech-to-English-text-translation (S2ETT), spoken language identification, and Voice activity detection on 680k hours of audio Radford et al. (2022). SeamlessM4Tv2-Large is a multilingual multimodal translation model Seamless Communication et al. (2023). It combines a text-to-text translation (T2TT) model pre-trained on 105 languages and a Conformer speech encoder pre-trained with wav2vec-BERT 2.0 on 4.5M hours of audio. The model is then trained on T2TT, S2TT, ASR, and knowledge distillation (KD) objectives. The authors perform KD from the text encoder to the speech encoder by minimizing the KL-divergence between the decoder’s token output distributions on bi-modal speech and text data.
Multilingual SLU. The evaluation of multilingual SLU is constrained to a limited set of tasks. SLU has been significantly shaped by task-oriented dialogue (ToD), with datasets created focusing on ToD-specific tasks such as intent classification and slot filling. These tasks frequently require only basic semantic understanding, often reducing to merely detecting specific keywords. Additionally, commonly used utterance-level SLU tasks like language identification (LID) and sentiment classification do not assess content-based understanding but instead rely on phonetic or prosodic features of speech encodings. As a result, the majority of SLU datasets are predominantly in English. Multilingual exceptions are otherwise limited. The Minds14 dataset for intent classification only includes 14 high-resource languages Gerz et al. (2021). SpeechTaxi offers spoken topical classification for Bible verses in 28 diverse languages, which however do not adequately represent real-world domains Keller and Glavaš (2024).
Concurrent Work. Costa-jussà et al. (2024) concurrently released the multilingual SLU benchmark 2M-Belebele for 74 languages. In this dataset, the authors first extend Flores by ca. 20% by incorporating human recordings for both sentences that are part of Flores but were missing in Fleurs, as well as the questions and answers from Belebele. Costa-jussà et al. (2024) find that a CS-based approach on transcribing speech with Whisper-v3-Large and subsequently prompting Llama 3 70B with the transcription trails prompting the LLM on clean text by ca. 8 percentage points (pp).
@@ 3 Fleurs-SLU
We postprocess datasets that are based on Flores to create ‘Fleurs-SLU’, a massively multilingual SLU benchmark for speech classification and multiple-choice QA from spoken paragraphs.
@@ 3.1 Core Datasets
Flores. The authors professionally translated 3,001 sentences of English Wikipedia paragraphs to evaluate machine translation Team et al. (2022).222Flores comprises 3,001 sentences divided into dev (997 sentences), devtest (1,012 sentences), and test (992 sentences) sets. The authors did not release the test set.
Fleurs. Fleurs comprises 2.3 spoken utterances, on average, per sentence from the dev and devtest splits of 102 languages in Flores Conneau et al. (2022).333Note that almost all languages part of Fleurs however are missing a few hundred sentences from Flores. Fleurs is used to evaluate multilingual ASR, LID, as well as speech-to-text and text-to-speech translation in all language directions.
SIB-200. Adelani et al. (2024) refined the topical metadata annotations of sentences in the dev and devtest splits of Flores into 7 categories.444The categories are ”science/technology”, ”travel”, ”politics”, ”sports”, ”health”, ”entertainment”, ”geography”. The resulting SIB-200 is a natural language understanding benchmark for 205 language variants.
Belebele. Belebele is a multiple-choice reading comprehension benchmark for 122 languages Bandarkar et al. (2024). The authors reconstruct paragraphs from Flores sentences. They generate 1–2 questions per English paragraph, which are professionally translated into 121 languages. Belebele comprises 900 questions spanning 488 passages.
@@ 3.2 Benchmark Creation
1) Merging Fleurs & Flores. We begin by removing silent instances from Fleurs.555Counts of removed examples are listed in Appendix A.6. We first normalize the loudness to an RMS target level of . We next apply voice activity detection using Silero-VAD Team (2024).666https://github.com/snakers4/silero-vad Samples are deemed silent if speech is detected in less than of their duration.777The Silero-VAD pipeline also frequently removes inaudibly noisy samples as a side effect. Lastly, we verify our approach on 50 randomly sampled predicted silent instances. We find only one borderline misclassified example, which is very noisy but still comprehensible. We then conservatively merge Fleurs and Flores by matching instances first on exact string match and then by Levenshtein distance of 3 on normalized strings.888Normalized strings remove characters that are Unicode punctuation codepoints.
2a) SIB-Fleurs. For each language, we pool instances from the training, validation, and test splits of SIB-200 Adelani et al. (2024) and align the data with our merged Fleurs-Flores dataset with the same string alignment procedure as before. The data is then regrouped into the training, validation, and test splits of the original Fleurs dataset. This segmentation ensures compatibility for speech models that may be trained on ASR using the training set of Fleurs prior to SIB-Fleurs evaluation. This also ensures that the speakers in the training set are different from those in the validation and test sets. Table 2 lists aggregated statistics on the instance- and the utterance-level for SIB-Fleurs. Appendix A.2 provides a list of samples by split per language.
| Task | Utterance |
@@
2b) Belebele-Fleurs. We merge our Fleurs-Flores sentences with Belebele paragraphs by intersecting the URLs of the texts. We discard all paragraphs that are not complete in Fleurs-Flores. We verify our reconstructed paragraphs against the original Belebele by ensuring that the Levenshtein distance for strings with removed punctuation is negligibly small (less than 3 characters). Table 2 provides a detailed summary statistics on both the paragraph- and the sentence-level for Belebele-Fleurs, while Appendix A.3 lists the samples per language.
Task Multiple-Choice QA Spoken Paragraphs Languages 92 Answer Choices 4 Questions per paragraph 1.8 (1.8) Sentences per paragraph 3.6 (4.0) Utterances per sentence 2.0 (2.0) Duration per utterance (s) 12.6 (12.0) Duration per paragraph (s) 47.0 (43.4) Total audio (hr) 944 Samples (Paragraphs) 709 (771)
@@ 4.1 Tasks and Languages
Belebele-Fleurs. We train and validate models on the English training and dev sets of Bandarkar et al. (2024), respectively. We jointly embed the paragraph, question, and choices with text encoders. We then average the token encodings of each choice and project the choice embedding via head to a logit . We minimize the cross-entropy of the concatenated choice logits to the label choice.999We do not evaluate speech LLMs as existing models are based on Whisper-v3 which is limited to 30 seconds of audio input. We leave such evaluation to future work.
@@ 4.2 Cross-Lingual Transfer Setups
We experiment on two commonly used cross-lingual transfer (xlt) paradigms. Zero-shot cross-lingual transfer (zs-xlt) and Translate-Test (TTest) allow us to evaluate xlt without requiring additional annotation for any target language. In zs-xlt, we first train a multilingual model on source-language data and then directly run inference on target-language instances. In TTest, the model is first fine-tuned on labeled source-language data. At test time, the target-language examples are translated to the source language prior to inference, which enables xlt with monolingual LLMs.
@@
Speech Classification. We evaluate state-of-the-art mSE for speech classification (cf. §2). We include MMS-1b without fine-tuning (‘MMS-1b’), with ASR fine-tuning on Fleurs (‘MMS-1b-Fleurs’), and with ASR fine-tuning on multilingual datasets (‘MMS-1b-all’). For ASR fine-tuning, Pratap et al. (2024) train language adapters and language-specific decoding heads while keeping all other parameters frozen. For task fine-tuning, we freeze the adapters to facilitate zs-xlt. We further evaluate mHubert and the speech encoders of both Whisper-v3-Large and SeamlessM4Tv2-Large.
@@
Cascading & Text. Cascaded systems (CS) perform xlt in two steps. CS first transcribe speech into text with an ASR model. The transcription is then fed into an text encoder paired with a classification head. We transcribe speech into the target language and into English with SeamlessM4Tv2-Large and Whisper-v3-Large. Note that speech-to-text-translation into English (S2ETT) with CS is equivalent to TTest as merely monolingual LLMs are needed for xlt. We further evaluate on the original text of SIB and Bebelebe, which we translate to English with SeamlessM4Tv2-Large.101010We use the gold Q&A for the respective language (i.e., English for ). Otherwise, we would need to interleave MT models, e.g., S2ETT of paragraphs with Whisper, but translating textual Q&A with SeamlessM4Tv2. We decode greedily in transcription, since more complex decoding strategies do not improve xlt performance Ebing and Glavaš (2024). We run our text-based and CS experiments on Roberta-Large Liu et al. (2019) as well as LLM2Vec and NLLB-LLM2Vec that are sequence encoders based on Llama 3 8B AI@Meta (2024). LLM2Vec BehnamGhader et al. (2024) enables bidirectional attention and continues training on both masked next-token prediction and SimCSE Gao et al. (2021). NLLB-LLM2Vec fuses LLM2Vec with the encoder of the NLLB translation model to extend its NLU capabilities to 200 languages Schmidt et al. (2024).
4.3 Training Setup

### v3

@@ Abstract
Spoken language understanding (SLU) is indispensable for half of all living languages that lack a formal writing system. Unlike for high-resource languages, for these languages, we cannot offload semantic understanding of speech to the cascade of automatic speech recognition (ASR) and text-based large language models (LLMs). Even if low-resource languages possess a writing system, ASR for these languages remains unreliable due to limited bimodal speech and text training data. Nonetheless, the evaluation of multilingual SLU is limited to shallow tasks such as intent classification or language identification. This is why we present Fleurs-SLU, a multilingual SLU benchmark that encompasses (i) 692 hours of speech for topical utterance classification in 102 languages and (ii) multiple-choice question answering via listening comprehension spanning 944 hours of speech across 92 languages. We extensively evaluate end-to-end speech classification models, cascaded systems that combine speech-to-text transcription with subsequent LLM-based classification, and multimodal speech-LLMs on Fleurs-SLU. Our results show that cascaded systems are more robust in multilingual SLU, though well-pretrained speech encoders can perform competitively in topical speech classification. Closed-source speech-LLMs match or surpass the performance of cascaded systems. We observe a strong correlation between robust multilingual ASR, effective speech-to-text translation, and strong multilingual SLU, indicating mutual benefits between acoustic and semantic speech representations. 111The datasets & per-language results are available on Huggingface: SIB-Fleurs; Belebele-Fleurs.
@@ 1 Introduction
Only about half of the world’s living languages possess a formal writing system, underscoring the need for massively multilingual speech technology (Ethnologue, 2017). Spoken language understanding (SLU) is hence a critical feature of inclusive technology for these languages, which cannot rely on combining automatic speech recognition (ASR) with language models to offload semantic speech understanding. However, the datasets currently available for evaluating SLU in truly relevant languages suffer from significant limitations. For instance, the Minds14 benchmark assesses SLU on 14 exclusively high-resource languages for intent classification, a task that often requires only shallow semantic processing, such as detecting specific keywords (Gerz et al., 2021). Likewise, while SpeechTaxi focuses on spoken topical classification of Bible verses in 28 diverse languages, SLU of Bible verse arguably does not reflect real-world usage (Keller & Glavaš, 2024).
@@
Moreover, even when a low-resource language has a writing system, multilingual ASR often struggles to reliably transcribe that language due to the limited availability of bimodal speech and text data. Pre-trained on the self-supervised wav2vec-BERT objective (Baevski et al., 2020), these models embed acoustic features that capture phonetic rather than semantic information (Choi et al., 2024). Conversely, multilingual speech models that excel in semantic encoding tend to be more robust in ASR tasks by, e.g., leveraging context-related cues or cross-lingual similarities such as cognates or similar syntactic structures (cf. §5.1). This underscores that robust SLU is instrumental for genuinely inclusive speech technology.
@@
To address the shortcomings of multilingual SLU evaluation of prior benchmarks, we compile Fleurs-SLU by realigning datasets derived from Flores (Team et al., 2022), a machine-translation benchmark with parallel sentences for over 200 languages. We first filter out silent instances from Fleurs (Conneau et al., 2022) and then map the remaining data back to Flores. The resulting dataset is then merged with SIB-200 (Adelani et al., 2024) and Belebele (Bandarkar et al., 2023), that both are based on Flores, to create a topical utterance classification benchmark for 102 languages, and dataset for multiple-choice question answering (QA) on spoken paragraphs for 92 languages, respectively.
@@
Contributions. 1) To the best of our knowledge, we are the first to present a multilingual SLU benchmark that spans over 100 languages and enables end-to-end SLU evaluation of multilingual speech encoders (mSE). SIB-Fleurs supports utterance classification for 7 topics with 692 hours of speech spanning 102 languages, over 70 more additional languages than existing benchmarks. Belebele-Fleurs provides textual multiple-choice QA from spoken paragraphs, totaling 944 hours of speech across 92 languages and exceeding concurrent work by 18 languages (Costa-jussà et al., 2024). 2) We extensively benchmark state-of-the-art speech models on SIB-Fleurs and Belebele-Fleurs. Like prior work, we observe that Cascaded Systems (CS) remain more robust than mSE, enabling SLU competitive with text-based language understanding (Keller & Glavaš, 2024). We additionally benchmark state-of-the-art speech-LLMs on multilingual SLU and find that often perform on par or better than CS. We are the first to show that mSE pre-trained on language understanding objectives can yield performance competitive to that of CS on speech classification. We further isolate utterance quality as an important factor in multilingual SLU and show that, in zero-shot cross-lingual transfer (zs-xlt), worse utterance quality can substantially deteriorate transfer performance. 3) We empirically demonstrate that strong multilingual SLU coincides with much more robust multilingual ASR and higher-quality speech-to-English-text translation (S2ETT). This suggests that the pre-training of multilingual speech models may take SLU into account to make multilingual ASR more robust.
@@ 2 Related Work
Modern speech representation models are pre-trained with self-supervised objectives that are optionally followed-up by ASR and speech-to-text translation (S2TT) training. Among these models, mHubert employs the self-supervised wav2vec 2.0 objective on 90K hours of speech across 147 languages by predicting pseudo-labels derived from clustering raw speech features (Boito et al., 2024). MMS-1b is also pretrained with wav2vec 2.0 on a large corpus of 491K hours of speech spanning 1,406 languages (Pratap et al., 2024). Whisper-v3 is a Transformer encoder-decoder that has been multi-task pre-trained on multilingual ASR, speech-to-English-text-translation (S2ETT), spoken language identification, and Voice activity detection on 680k hours of audio (Radford et al., 2022). SeamlessM4Tv2-Large is a multilingual multimodal translation model (Seamless Communication et al., 2023). It combines a text-to-text translation (T2TT) model pre-trained on 105 languages and a Conformer speech encoder pre-trained with w2v-BERT 2.0 on 4.5M hours of audio. The model is then trained on T2TT, S2TT, ASR, and knowledge distillation (KD) objectives. The authors perform KD from the text encoder to the speech encoder by minimizing the KL-divergence between the decoder’s token output distributions on bi-modal speech and text data.
Multilingual SLU. The evaluation of multilingual SLU is constrained to a limited set of tasks. SLU has been significantly shaped by task-oriented dialogue (ToD), with datasets created focusing on ToD-specific tasks such as intent classification and slot filling. These tasks frequently require only basic semantic understanding, often reducing to merely detecting specific keywords. Additionally, commonly used utterance-level SLU tasks like language identification (LID) and sentiment classification do not assess content-based understanding but instead rely on phonetic or prosodic features of speech encodings. As a result, the majority of SLU datasets are predominantly in English. Multilingual exceptions are otherwise limited. The Minds14 dataset for intent classification only includes 14 high-resource languages (Gerz et al., 2021). SpeechTaxi offers spoken topical classification for Bible verses in 28 diverse languages, which however do not adequately represent real-world domains (Keller & Glavaš, 2024).
Concurrent Work. Costa-jussà et al. (2024) concurrently released the multilingual SLU benchmark 2M-Belebele for 74 languages. In this dataset, the authors first extend Fleurs by approx. 20% by incorporating human recordings for sentences that are part of Flores but were missing in Fleurs, as well as the questions and answers from Belebele. Costa-jussà et al. (2024) find that a CS-based approach on transcribing speech with Whisper-v3-Large and subsequently prompting Llama 3 70B with the transcription trails prompting the LLM on clean text by approx. 8 percentage points (pp).
@@ 3 Fleurs-SLU
We create Fleurs-SLU, a massively multilingual SLU benchmark for speech classification and multiple-choice QA from spoken paragraphs, from datasets that are based on Flores.
@@ 3.1 Core Datasets
Flores consists of professionally translated 3,001 sentences of English Wikipedia paragraphs to evaluate machine translation (Team et al., 2022).222Flores comprises 3,001 sentences divided into dev (997 sentences), devtest (1,012 sentences), and test (992 sentences) sets. The authors did not release the test set. Fleurs comprises 2.3 spoken utterances, on average, per sentence from the dev and devtest splits of 102 languages in Flores (Conneau et al., 2022).333Almost all languages part of Fleurs however are missing a few hundred sentences from Flores. Fleurs is used to evaluate multilingual ASR, LID, as well as speech-to-text and text-to-speech tanslation in all language directions. SIB-200 refined the topical metadata annotations of sentences in the dev and devtest splits of Flores into 7 categories (Adelani et al., 2024).444”science/technology”, ”travel”, ”politics”, ”sports”, ”health”, ”entertainment”, ”geography”. The resulting SIB-200 is a topical classification benchmark for 205 language variants. Belebele is a multiple-choice reading comprehension benchmark for 122 languages (Bandarkar et al., 2024). The authors reconstruct paragraphs from Flores sentences. They generate 1-2 questions per English paragraph, which are professionally translated into 121 languages. Belebele comprises 900 questions that span across 488 passages.
@@ 3.2 Benchmark Creation
1) Merging Fleurs & Flores. We begin by removing silent instances from Fleurs.555Counts of removed examples are listed in Appendix A.6. We first normalize the loudness to a target RMS level of dB. We next apply voice activity detection using Silero-VAD (Team, 2024).666https://github.com/snakers4/silero-vad Samples are deemed silent if speech is detected in less than of their duration.777The Silero-VAD pipeline also frequently removes inaudibly noisy samples as a side effect. Lastly, we verify our approach on 50 randomly sampled predicted silent instances. We find only one borderline misclassified example, which is noisy but comprehensible. We then conservatively merge Fleurs and Flores by matching instances first on exact string match and then by Levenshtein distance of 3 on normalized strings.888Normalized strings remove characters that are Unicode punctuation codepoints.
2a) SIB-Fleurs. For each language, we pool instances from the training, validation, and test splits of SIB-200 (Adelani et al., 2024) and align the data with our merged Fleurs-Flores dataset with the same string alignment procedure as before. The data is then regrouped into the training, validation, and test splits of the original Fleurs dataset. This segmentation ensures compatibility for speech models that may be trained on ASR using the training set of Fleurs prior to SIB-Fleurs evaluation. This also ensures that the speakers in the training set are different from those in the validation and test sets. Table 2 lists aggregated statistics on the instance- and the utterance-level for SIB-Fleurs. Appendix A.2 provides a list of samples by split per language.
Utterance | |
@@
Multiple-Choice QA | |
Spoken Paragraphs | |
Languages | 92 |
Answer Choices | 4 |
Questions per paragraph | 1.8 (1.8) |
Sentences per paragraph | 3.6 (4.0) |
Utterances per sentence | 2.0 (2.0) |
Duration per utterance (s) | 12.6 (12.0) |
Duration per paragraph (s) | 47.0 (43.4) |
Total audio (hr) | 944 |
Samples (Paragraphs) | 709 (771) |
2b) Belebele-Fleurs. We merge our Fleurs-Flores sentences with Belebele paragraphs by intersecting the URLs of the texts. We discard all paragraphs that are not complete in Fleurs-Flores. We verify our reconstructed paragraphs against the original Belebele by ensuring that the Levenshtein distance for strings with removed punctuation is negligibly small (less than 3 characters). Table 2 provides a summary statistics on both the paragraph- and the sentence-level for Belebele-Fleurs, while Appendix A.3 lists the samples per language.
@@ 4.1 Tasks and Languages
Belebele-Fleurs. We train and validate CS models on the English training and dev sets of Belebele (Bandarkar et al., 2024), respectively. We jointly embed the paragraph, question, and choices with text encoders. We then average the token encodings of each choice and project the choice embedding via head to a logit . We minimize the cross-entropy of the concatenated choice logits to the label choice.999We do not evaluate speech LLMs as existing models are based on Whisper-v3 which is limited to 30 seconds of audio input. We leave such evaluation to future work.
@@ 4.2 Cross-Lingual Transfer Setups
We experiment on two commonly used cross-lingual transfer (xlt) paradigms. Zero-shot cross-lingual transfer (zs-xlt) and Translate-Test (TTest) allow us to evaluate xlt without requiring additional annotation for any target language. In zs-xlt, we first train a multilingual model on the English source-language data (cf. §4.1) and then directly run inference on the target-language test instances. In TTest, the model is also first fine-tuned on labeled English source-language data. At test time, the target-language examples are translated to the source language prior to inference, which enables xlt with monolingual LLMs. We additionally evaluate zero-shot prompting of speech-LLMs. Here, we provide the context of the task in English and all relevant input in-language in the respective modality, i.e. text or utterances, to the speech-LLMs (cf. §A.1).
@@
Speech Classification. We evaluate state-of-the-art mSE for speech classification (cf. §2). We include MMS-1B without fine-tuning (‘MMS-1B’), with ASR fine-tuning on Fleurs (‘MMS-1B-Fleurs’), and with ASR fine-tuning on multilingual datasets (‘MMS-1B-all’), allowing us to analyze the impact of ASR fine-tuning on cross-lingual SLU. For ASR fine-tuning, Pratap et al. (2024) train language adapters and language-specific decoding heads while keeping all other parameters frozen. For task fine-tuning, we freeze the adapters to facilitate zs-xlt. We further evaluate mHubert and the speech encoders of both Whisper-v3-Large and SeamlessM4Tv2-Large (cf. §2).
@@
Cascading & Text. Cascaded systems (CS) perform xlt in two steps: (i) an ASR model transcribes speech into text, and (ii) a text encoder processes the transcription via a classification head. We consider two transcription targets: the target language (in-language) and English (speech-to-English translation, S2ETT). S2ETT corresponds to TTest, enabling xlt with monolingual LLMs. We use SeamlessM4Tv2-Large and Whisper-v3-Large as ASR backends. For languages not supported by the model, we manually select the closest available language for in-language transcription. All decoding is greedy, as more complex decoding strategies do not improve xlt performance (Ebing & Glavaš, 2024). We additionally evaluate on the original text (Text) of SIB-200 and Belebele. For TTest, we translate this text to English using SeamlessM4Tv2-Large.101010We use the gold Q&A in English for TTest. Otherwise, MT models would need to be combined, e.g., S2ETT of paragraphs with Whisper, and translation of textual Q&A with SeamlessM4Tv2. For classification, we evaluate three text encoders: Roberta-Large (Liu et al., 2019), LLM2Vec (BehnamGhader et al., 2024), and NLLB-LLM2Vec (Schmidt et al., 2024). LLM2Vec is a sequence encoder based on Llama 3 8B (AI@Meta, 2024), trained with bidirectional attention on masked next-token prediction and SimCSE (Gao et al., 2021). NLLB-LLM2Vec extends LLM2Vec with the encoder of the NLLB translation model, covering 200+ languages for robust multilingual NLU (Schmidt et al., 2024).
Speech-LLMs. We further evaluate two multimodal speech-LLMs, Qwen 2.5 7B-Omni (QwenOmni) and Gemini 2.0 Flash111111https://cloud.google.com/vertex-ai/generative-ai/docs/models/gemini/2-0-flash, in a zero-shot prompting setup.121212Note that the audio encoder of Qwen 2.5 7B-Omni is initialized with Whisper-v3-Large prior to further cross-modal instruction tuning. For each task, we provide an English task description and supply the task-specific input (e.g., question, paragraph, or utterance) in the target language, using the appropriate modality (text or speech). To ensure consistent audio quality across languages, we normalize all utterances to a root mean square (RMS) level of 0.07. For Belebele-Fleurs, where inputs consist of multiple utterances, we normalize the concatenated sequence. Appendix §A.1 details the prompt formats and input preprocessing used for each task and modality.
4.3 Further Details